Each month, a Filipino household generates an average of 30 pounds of garbage for recycling. Assume that the distribution is normal with a standard deviation of 2 pounds. If a household is selected at random, find the probability of its generating more than 34 pounds per month. *




1
Expert's answer
2022-03-01T17:44:52-0500

"P(X>34)=P(Z>\frac{34-30}{2})=P(Z>2)=1-P(Z<2)=0.0228."